OpenStudy (anonymous):

need help with some exponents, 7x^(3/2)-21x^(1/2)+14x^(-1/2) Any help is appreciated.

OpenStudy (anonymous):

\[7x ^{\frac{ 3 }{ 2 }}-21x ^{\frac{ 1 }{ 2 }}-14x ^{\frac{ -1 }{ 2 }}\]

OpenStudy (anonymous):

what do you need to be done here, what is the question saying @freshmanqs

OpenStudy (amoodarya):

factor x^(-1/2) to simplify

OpenStudy (amoodarya):

and factor 7

OpenStudy (amoodarya):

another way to simplify is substitute a=x^(-1/2)

OpenStudy (amoodarya):

\[7x^{\frac{-1}{2}} (x^2-3x+2)=\\7x^{\frac{-1}{2}} (x-1)(x-2)\]